# Low bit-flip rate probabilistic error cancellation

### v1

@@ Abstract
Noise remains one of the most significant challenges in the development of reliable and scalable quantum processors. While quantum error correction and mitigation techniques offer potential solutions, they are often limited by the substantial hardware overhead required. To address this, tailored approaches that exploit specific hardware characteristics have emerged. In quantum computing architectures utilizing cat-qubits, the inherent exponential suppression of bit-flip errors can significantly reduce the qubit count needed for effective error correction. In this work, we explore how the unique noise bias of cat-qubits can be harnessed to enhance error mitigation efficiency. Specifically, we demonstrate that the sampling cost associated with probabilistic error cancellation (PEC) methods can be substantially lowered when applied to circuits built on cat-qubits, provided the gates used preserve the noise bias. Our error mitigation scheme is benchmarked across various quantum machine learning circuits, showcasing its practical advantages.
I Introduction
@@
Quantum error correction techniques aim to revert the computational errors which unavoidably arise during the execution of quantum circuits. However, the physical qubit count overhead they impose make them impractical for smaller machines.
Quantum error mitigation restricts the task of handling errors to improving the accuracy of the expectation values obtained by sampling the output from quantum circuits. In doing so, the cost of handling errors shifts from a space overhead (the qubit count required to correct errors) to a time overhead (the number of times one needs to run quantum circuits to mitigate errors). In essence, error mitigation aims to minimize the space overhead in an attempt to make the smallest possible quantum devices deliver useful computational power.
In this context, Probabilistic Error Cancellation (PEC) is an error mitigation technique that relies on the possibility to approximate the expectation value of an observable measured on a quantum state prepared by a noiseless circuit as the average value of measurement results taken over a quasi-probabilistic distribution of noisy quantum circuits (see zne-pec for the original techniques and, e.g. practical-pec ; optimal-pec ; limits-pec ; ECBY21hybrid for improvements and guarantees). This allows to replace the execution of a noiseless circuit with that of noisy but easily implementable ones, while yielding an unbiased estimator for the expectation value of observables on the noiseless circuits.
This method has been put to work for numerous use cases, for example for superconducting quantum architecture with crosstalk BMKT23probabilistic , or in the mitigation of Markovian noise EBL18practical , or in a setting in which one learns the noise associated to a given system SQCB21learning .
Thanks to its appealing feature of being hardware and algorithm agnostic, PEC is easy to use. Yet, this also means that PEC does not exploit knowledge of error propagation within the quantum circuits of interest in order to perform more efficient simulations. One might thus ask, following (CBBE23quantum, , Open Question 3), whether PEC can be further optimized if one is willing to specialize it to a given hardware or algorithm.
In this paper, we provide a positive answer to the previous open question. We do so by considering cat-qubit architectures QCwithCats . Using a biased noise error model, and with the knowledge of the propagation of errors within bias-preserving circuits, we introduce a variant of PEC that outperforms standard PEC in terms of sampling cost overhead for achieving a given target variance for the expectation value of an observable measured at the output of the noiseless circuit. We derive general theoretical performance guarantees, provide numerical explorations of our method and describe their impact for specific machine learning applications.
@@
PEC operates by sampling noisy circuits to construct an approximation of an ideal noiseless circuit given as input. When such ideal circuit is provided as a sequence of layered gates, PEC constructs a collection of noisy circuits from each layer by independently adding extra gates that mitigate the effect of the noise on average. This unfortunately results in a quantum sampling cost that increases exponentially with the number of layers.
To improve over such a situation, we seek to reduce the base and exponent of the exponential quantum sampling cost that PEC incur. We do so by grouping the added gates together by commuting them through the various noisy layers of the circuit. They are then combined into a single additional layer. This is effectively performing PEC through block sampling. We thus refer to this method as Block-PEC. We then derive theoretical guarantees to obtain performance improvements. It shows that Block-PEC trades part of the quantum sampling cost for a classical pre-processing step which calculates quasi-probabilities. This results in a quantum sampling exponential in the qubit count of the circuit rather than exponential in the qubit count and the circuit depth as is the case for conventional PEC approach.
We then show that cat-qubit architectures — architectures for which bit-flips are exponentially suppressed with a parameter that can be tuned experimentally QCwithCats — via their 1 and 2-qubit bias-preserving gates, mapping phase flips to phase flips, naturally satisfy the conditions required for efficient Block-PEC. It thus provides the first quantum error mitigation technique which specifically targets such architectures. We then relax the requirements of Block-PEC by considering circuits which are bias-preserving on a given subspace of the full Hilbert space accessible to the computer.
Finally, we explore the practical performance of this technique through numerical simulations using small size circuits corresponding to a variety of bias-preserving circuits at the core of machine learning algorithms and computational finance applications of quantum computing. This is done using various biased noise model assumptions (i.e. with and without correlated errors for 2-qubit gates). We find a clear improvement for parameterized quantum circuits formed by matchgates, and in particular Reconfigurable Beam Splitter (RBS) gates, which are widely used in quantum machine learning algorithms for NISQ computers, see e.g. unary-option-pricing ; rbs-pde ; low-depth-qae ; qdeep-hedging ; qvt .
@@
PEC is a well established error mitigation technique available in various open source packages (see e.g. mitiq ). From optimal-pec , we know that standard PEC is optimal (in terms of sampling cost) when applied on one $n$-qubit layer subjected to dephasing noise. Moreover, it has been established that using a continuous set of ideal gates, containing all $Z$-axis rotations for arbitrary angles, would not help lower the cost of PEC, as explained in (optimal-pec, , Sec. 5).
Previous works have nonetheless provided improvements to PEC by optimizing the sampling cost taking into account the gates within the past light-cone of a given observable, as only those contribute to the errors affecting the measured expectation value local-qem .
@@
After an overview of the mathematical formalism of probabilistic error cancellation (Section II), we introduce Block-PEC, and we outline the conditions under which it can be implemented (Section III). We then provide performance guarantees for Block-PEC (Section IV), followed by a numerical analysis of the implementation of Block-PEC on concrete applications (Section V). We conclude in Section VI, focusing on the prospects of this approach, and placing it in the context of hardware-specific error mitigation techniques.
II Preliminaries
II.1 Quantum channels
@@
Definition II.1 (Noisy unitary channel).
@@
II.2 cat-qubits and dephasing noise
cat-qubits are superconducting qubits which can be physically implemented in such a way that bit flip errors are exponentially suppressed, at the cost of a linear increase of the phase flip error rate (see QCwithCats for a detailed presentation of cat-qubits). In other words, cat-qubits are associated with a biased noise model, heavily biased towards phase flip errors.
There is experimental evidence that phase flips errors are drastically rarer than bit flip errors ($10^{8}$ times rarer in the experiments in 100secs-bitflip , for example). In the present work, we choose to ignore bit flips altogether (we further explain this choice in Appendix B). To make our results more concrete, we make use of two simplified dephasing noise models. The first one inserts correlated phase flips after each gate on the qubits it has been acting upon. More precisely, we chose to consider that after a gate acting on the set of $n$ qubits $A$, the following channel acts upon these qubits:
@@
so that $Z_{\emptyset}=I^{\otimes n}$. Such correlations have been observed in cat-qubit architectures for 1 and 2-qubit gates (see e.g. QCwithCats ). The second model is a more drastic approximation but yields simpler analytic calculations. It consists of applying single qubit dephasing channels on all the qubits a gate is acting upon. In effect, for a gate acting on the set of qubits $A$, the channel applied after the gate is
@@
II.3 Overview of Probabilistic Error Cancellation
To compensate the effect of noise, error mitigation techniques such as PEC intend to apply the inverse of the noise channel to the affected quantum systems. While it is not a physically realizable operation (when the noise channel is not unitary), it is nonetheless well defined mathematically.
Indeed (see zne-pec ; practical-pec and also Appendix A), for large classes of noise models, one can rewrite the channel $\mathsf{U}$ corresponding to a noiseless $d$-layered circuit as a quasi-probabilistic distribution of noisy circuits practical-pec ; zne-pec :
@@ II.3 Overview of Probabilistic Error Cancellation
where $\sum_{\mathsf{V}_{l}}\alpha_{l}(\mathsf{V}_{l})=1$, and $n_{l}$ is the number of qubits that the $l$-th layer $U_{l}$ acts non-trivially on.
Under the assumption that the noisy unitary channels $\widetilde{\mathsf{B}_{l}}=\mathsf{N}_{l}\circ\mathsf{V}_{l}\circ\mathsf{U}_{l}$ correspond to physically implementable gates, PEC determines the expectation value of the noiseless circuit $U$ by randomly sampling a noisy gate $\widetilde{\mathsf{B}_{l}}$ for each layer, and executing the circuit corresponding to $\widetilde{\mathsf{B}_{d}}\cdots\widetilde{\mathsf{B}_{1}}$.
@@
In a conventional use of PEC, the sampling cost $\gamma_{\text{total}}$ is in $O(\gamma^{nd})$, where $n$ is the qubit count, $d$ is the circuit depth, and $\gamma\simeq 1+kp+O(p^{2})$ for some constant $k$ determined by the noise model (for dephasing noise, $k=2$), with $1$-qubit error rate $p$. This makes PEC a reasonable mitigation strategy in the presence of weak gate noise optimal-pec .
II.4 Inverting Pauli noise
@@
Furthermore, if the non trivial Pauli-channels in $\mathsf{A}$ correspond to Pauli operators that form a group, so do the non trivial terms in the quasi-probabilistic distribution of $N^{-1}$.
$N^{-1}$ can be approximated by a channel of the form $(1+p)\mathsf{I}-p\mathsf{A}$.
@@
$N^{-1}$ $=\left((1-p)\mathsf{I}+p\mathsf{A}\right)^{-1}$ (10)
@@
The first and second claim of the theorem follow from the group structure of $\mathbb{Q}$ which guarantee that each term in the expansion of $A^{i}=\left(\sum_{\mathsf{Q}\in\mathbb{Q}}p_{\mathsf{Q}}Q\right)^{i}$ is in $\mathbb{Q}$.
@@
III Block Probabilistic Error Cancellation
The standard PEC procedure outlined in Section II.3 is oblivious to the structure of the implementable gate set and noise model save for the actual quasi-probabilistic decomposition of each noisy gate. In this section, we go beyond this generic approach by using a property of the implementable gate set that we call Pauli-$Z$ compatibility to modify the Monte Carlo sampling performed in the PEC procedure.
@@ III.1 Pauli-$Z$ compatibility
Definition III.1 (Pauli-$Z$ compatibility).
@@
A Pauli-$Z$ string is a $n$-fold tensor product of single qubit $I$ and $Z$. Whenever $\mathbb{N}$ is the set of all convex combinations of Pauli-$Z$ strings, we say that $(\mathbb{U},\mathbb{V})$ is Pauli-$Z$ compatible.
For convenience, we will say that a gate is Pauli-$Z$ compatible if is in a set $\mathbb{U}$ such that $(\mathbb{U},\mathbb{V})$ is compatible, with $\mathbb{V}$ being the set of all Pauli-$Z$ strings. A circuit is Pauli-$Z$ compatible if it is generated from a Pauli-$Z$ compatible gate set.
@@
III.2 Block probabilistic error cancellation
Consider a layered circuit $\mathsf{U}=\mathsf{U}_{d}\circ\cdots\circ\mathsf{U}_{1}$ with layers $\mathsf{U}_{l}\in\mathbb{U}$, each associated to their own noise channel $\mathsf{N}_{l}$, forming a non-uniform noise model.
Each layer $\mathsf{U}_{l}$ can be decomposed into a quasi-probabilistic distribution of layers obtained by applying a perfect additional control:
@@ III.2 Block probabilistic error cancellation
We can now commute all the additional controls $\mathsf{V}_{l}$ to the end of the circuit resulting into a quasi-probabilistic distribution corresponding to $\mathsf{U}$:
@@
III.3 Computing Block-PEC coefficients

### v2

@@ Abstract
Noise remains one of the most significant challenges in the development of reliable and scalable quantum processors. While quantum error correction and mitigation techniques offer potential solutions, they are often limited by the substantial overhead required. To address this, tailored approaches that exploit specific hardware characteristics have emerged. In quantum computing architectures utilizing cat-qubits, the inherent exponential suppression of bit-flip errors can significantly reduce the qubit count needed for effective error correction. In this work, we explore how the unique noise bias of cat-qubits can be harnessed to enhance error mitigation efficiency. Specifically, we demonstrate that the sampling cost associated with probabilistic error cancellation (PEC) methods can be exponentially reduced with the depth of the circuit when gates act on cat-qubits and preserve the noise bias. Similar results also hold for Clifford circuits and Pauli channels. Our error mitigation scheme is benchmarked across various quantum machine learning circuits, showcasing its practical advantages for near-term applications on cat-qubit architectures.
1 Introduction
@@
Quantum error correction techniques aim to undo computational errors that unavoidably arise during the execution of quantum circuits. However, the physical qubit count overhead they impose make them impractical for smaller machines. In contrast, quantum error mitigation techniques focus on improving the accuracy of expectation values obtained by sampling the output from quantum circuits, thereby shifting the cost of handling errors from a space overhead to a time overhead. Although the scaling of the overhead for error mitigation is not favorable for larger circuits, it is currently the best approach to address noise for small quantum devices and unlocking their computational power for near term applications.
One such technique is Probabilistic Error Cancellation (PEC), which approximates the expectation value of an observable measured on a quantum state prepared by a noiseless circuit as the average value of measurement results taken over a quasi-probabilistic distribution of noisy quantum circuits (see [32] for the original techniques and, e.g. [10, 30, 31, 12] for improvements and guarantees). By replacing the execution of a noiseless circuit with that of noisy but easily implementable ones, PEC yields an unbiased estimator for the expectation value of observables on the noiseless circuits.
Thanks to its appealing feature of being hardware and algorithm agnostic, PEC is easy to use and has been successfully applied to various use cases, including superconducting quantum architectures with crosstalk [34], mitigation of Markovian noise [11], and learning the noise associated with a given system [29].
Yet, this also means that PEC remains unaware of the error propagation structure within the quantum circuits of interest, potentially leading to suboptimal performance. This raises an important question, posed in [6, Open Question 3]: can PEC be further optimized by tailoring it to a specific hardware or algorithm? Among current hardware proposal, cat-qubits—superconducting qubits with highly biased noise, considered as promising candidates for building fault-tolerant architectures [16]—are of great interest. Can their unique properties be leveraged to reduce the overhead of error mitigation? More broadly, do cat-qubits offer advantages outside the fault-tolerant regime and into the NISQ regime, where error mitigation is crucial for unlocking the potential of near-term quantum devices?
@@
In this paper, we provide a positive answer to the previous open questions. We do so by introducing a variant of Probabilistic Error Cancellation (PEC) that outperforms the standard approach in terms of sampling cost overhead for achieving a given target variance for the expectation value of an observable measured at the output of the noiseless circuit.
Indeed, the overhead of PEC is due to the necessity to sample noisy circuits to construct an approximation of an ideal noiseless circuit given as input. More precisely, when the ideal circuit is provided as a sequence of layered gates, PEC constructs a collection of noisy circuits from each layer by independently adding extra gates that mitigate the effect of the noise on average. This results in a quantum sampling cost that increases exponentially with the number of layers.
To address this limitation, we propose a modified error mitigation procedure, which we refer to as Block-PEC. By grouping the added gates together and commuting them through the various noisy layers of the circuit, we can combine them into a single additional layer, effectively performing PEC through block sampling. This allows us to reduce the base and exponent of the exponential quantum sampling cost, trading part of the quantum sampling cost for a classical pre-processing step that calculates quasi-probabilities. As a result, Block-PEC achieves a quantum sampling cost exponential in the qubit count of the circuit, rather than exponential in both the qubit count and the circuit depth. We complement our theoretical analysis by giving performance guarantees under various biased noise models as well as numerical evalutation of the gain for several concrete circuit classes.
By considering circuits that are bias-preserving on a given subspace of the full Hilbert space, we can further relax the requirements of Block-PEC. This allows us to demonstrate that Block-PEC can be effectively applied to circuits with biased noise. Our numerical simulations using small-size circuits show that Block-PEC provides a clear improvement for parameterized quantum circuits formed by matchgates, particularly Reconfigurable Beam Splitter (RBS) gates, which are widely used in quantum machine learning algorithms for NISQ computers [25, 17, 15, 8, 7].
@@
PEC is a well established error mitigation technique available in various open source packages (see e.g. [21]). From [30], we know that standard PEC is optimal (in terms of sampling cost) when applied on one $n$-qubit layer subjected to dephasing noise. Moreover, it has been established that using a continuous set of ideal gates, containing all $Z$-axis rotations for arbitrary angles, would not help lower the cost of PEC, as explained in [30, Sec. 5].
Previous works have nonetheless provided improvements to PEC by optimizing the sampling cost taking into account the gates within the past light-cone of a given observable, as only those contribute to the errors affecting the measured expectation value [33].
@@
After an overview of the mathematical formalism of probabilistic error cancellation (Section 2), we introduce Block-PEC, and we outline the conditions under which it can be implemented (Section 3). We then provide performance guarantees for Block-PEC (Section 4), followed by a numerical analysis of the implementation of Block-PEC on concrete applications (Section 5). We conclude in Section 6, focusing on the prospects of this approach, and placing it in the context of hardware-specific error mitigation techniques.
2 Preliminaries
2.1 Quantum channels
@@
Definition 2.1 (Noisy unitary channel).
@@
Throughout the rest of the manuscript “tilded” channels refer to their noisy version.
@@
2.2 Cat-qubits and dephasing noise
Cat-qubits are superconducting qubits which can be physically implemented in such a way that bit flip errors are exponentially suppressed, at the cost of a linear increase of the phase flip error rate (see [16] for a detailed presentation of cat-qubits). In other words, cat-qubits are associated with a biased noise model, heavily biased towards phase flip errors.
There is experimental evidence that phase flips errors are drastically rarer than bit flip errors ($10^{8}$ times rarer in the experiments in [5], for example). In the present work, we choose to ignore bit flips altogether (we further discuss this choice in Appendix B). To make our results more concrete, we make use of two simplified dephasing noise models.
The first one inserts correlated phase flips immediately after each gate of the circuit. More precisely, we chose to consider that after a gate acting on the set of $n$ qubits $A$, the following channel acts upon these qubits:
@@ 2.2 Cat-qubits and dephasing noise
so that $Z_{\emptyset}=I^{\otimes n}$. Such correlations have been observed in cat-qubit architectures for 1 and 2-qubit gates (see e.g. [16]).
The second model is a more drastic approximation but yields simpler analytic calculations. It consists of applying single qubit dephasing channels on all the qubits a gate is acting upon. In effect, for a gate acting on the set of qubits $A$, the channel applied after the gate is
@@
2.3 Overview of Probabilistic Error Cancellation
To compensate the effect of noise, error mitigation techniques such as PEC intend to apply the inverse of the noise channel to the affected quantum systems. While it is not a physically realizable operation as soon as the noise channel is not unitary, it is nonetheless well defined mathematically.
Indeed (see [32, 10] and also Appendix A), for large classes of noise models, one can rewrite the channel $\mathsf{U}$ corresponding to a noiseless $d$-layered circuit as a quasi-probabilistic distribution of noisy circuits [10, 32]:
@@ 2.3 Overview of Probabilistic Error Cancellation
where $\sum_{\mathsf{V}_{l}}\alpha_{l}(\mathsf{V}_{l})=1$, and $n_{l}$ is the number of qubits that the $l$-th layer $\mathsf{U}_{l}$ acts non-trivially on.
Under the assumption that the noisy unitary channels $\widetilde{\mathsf{B}_{l}}=\mathsf{N}_{l}\circ\mathsf{V}_{l}\circ\mathsf{U}_{l}$ correspond to physically implementable gates, PEC determines the expectation value of the noiseless circuit $\mathsf{U}$ by randomly sampling a noisy gate $\widetilde{\mathsf{B}_{l}}$ for each layer, and executing the circuit corresponding to $\widetilde{\mathsf{B}_{d}}\cdots\widetilde{\mathsf{B}_{1}}$.
@@
In a conventional use of PEC, the sampling cost $\gamma_{\text{total}}$ is in $O(\gamma^{nd})$, where $n$ is the qubit count, $d$ is the circuit depth, and $\gamma\simeq 1+kp+O(p^{2})$ for some constant $k$ determined by the noise model (for dephasing noise, $k=2$), with $1$-qubit error rate $p$. This makes PEC a reasonable mitigation strategy in the presence of weak gate noise [30].
2.4 Inverting Pauli noise
@@
Furthermore, if the non trivial Pauli-channels in $\mathsf{A}$ correspond to Pauli operators that form a group, so do the non trivial terms in the quasi-probabilistic distribution of $\mathsf{N}^{-1}$.
$\mathsf{N}^{-1}$ can be approximated by a channel of the form $(1+p)\mathsf{I}-p\mathsf{A}$.
@@
$\mathsf{N}^{-1}$ $=\left((1-p)\mathsf{I}+p\mathsf{A}\right)^{-1}$ (10)
@@
The first and second claim of the theorem follow from the group structure of $\mathbb{Q}$ which guarantee that each term in the expansion of $\mathsf{A}^{i}=\left(\sum_{\mathsf{Q}\in\mathbb{Q}}p_{\mathsf{Q}}\mathsf{Q}\right)^{i}$ is in $\mathbb{Q}$.
@@
Note that this inversion assumes that the noise channel $\mathsf{N}$ is close to the identity. In accordance with this assumption, the gates $\widetilde{\mathsf{B}_{l}}$ that will be used to invert $\mathsf{N}$ will introduce a noise that will be of similar magnitude. This means, following Eq, 12, that this effect will be of order 2 and can thus be neglected in our computation.
3 Block Probabilistic Error Cancellation
The standard PEC procedure outlined in Section 2.3 is oblivious to the structure of the implementable gate set and noise model save for the actual quasi-probabilistic decomposition of each noisy gate. In this section, we go beyond this generic approach by using a property of the implementable gate set that we call Pauli-$Z$ compatibility to modify the Monte Carlo sampling performed in the PEC procedure.
@@ 3.1 Pauli-$Z$ compatibility
Definition 3.1 (Pauli-$Z$ compatibility).
@@
In such case, $(\mathbb{U},\mathbb{V})$ is said to be $\mathbb{N}$-compatible. When $\mathbb{N}$ is the set of convex combinations of unitary channels made of Pauli-$Z$ strings — i.e. $n$-fold tensor product of single qubit $I$ and $Z$ — we say that $(\mathbb{U},\mathbb{V})$ is Pauli-$Z$ compatible.
For convenience, we will say that a gate is Pauli-$Z$ compatible if is in a set $\mathbb{U}$ such that $(\mathbb{U},\mathbb{V})$ is Pauli-$Z$ compatible. A circuit is Pauli-$Z$ compatible if it is generated from a Pauli-$Z$ compatible gate set.
@@
Note that if one instead restricts $\mathbb{U}$ to a subset of the Clifford group such that it contains the Pauli group, then for $\mathbb{V}$ the Pauli group and $\mathbb{N}$ the set of Pauli channels, properties P1 to P3 above are satisfied so that it can be said that $(\mathbb{U},\mathbb{V})$ is Pauli-compatible.
@@
3.2 Block probabilistic error cancellation
Consider $(\mathbb{U},\mathbb{V})$ and $\mathbb{N}$ satisfying P1 to P3 above, and a layered circuit $\mathsf{U}=\mathsf{U}_{d}\circ\cdots\circ\mathsf{U}_{1}$ with layers $\mathsf{U}_{l}\in\mathbb{U}$, each associated to their own noise channel $\mathsf{N}_{l}$, forming a non-uniform noise model. Each layer $\mathsf{U}_{l}$ can be decomposed into a quasi-probabilistic distribution of layers obtained by applying a perfect additional control:
@@ 3.2 Block probabilistic error cancellation
We can now commute all the additional controls $\mathsf{V}_{l}$ to the end of the circuit. This results into a quasi-probabilistic distribution corresponding to $\mathsf{U}$:
@@
Equations 20 and 21 provide two insights in the structure of the advantage one can gain by applying Block-PEC over standard PEC. First, whenever it can be applied, it is never detrimental to the sampling overhead. Second, and most importantly, bigger gains will be obtained whenever, for a given $\mathsf{V}^{\prime}\in\mathbb{V}$, the sum of the contributing $\alpha_{\mathsf{U}_{1}}(\mathsf{V}_{1})\cdots\alpha_{\mathsf{U}_{d}}(\mathsf{V}_{d})$ terms to $\mathsf{V}^{\prime}$ in Equation 21 is small compared to the sum of their absolute values in Equation 20. This requires having terms with different signs, and as there are a number of terms that grows exponentially with $d$, one can expect up to an exponential gain in $d$ by using Block-PEC. We will show in Section 4 that this is the case for broad classes of circuits and realistic noise models. Section 5 explore the gain on end-to-end applications, some of which can run entirely on bias-preserving gates, adding to the interest of Block-PEC for cat-qubits.
We will show in Section 4 that this is the case for broad classes of circuits and realistic noise models with bias-preserving gates. Section 5 explores the gain on end-to-end applications, some of which can run entirely on bias-preserving gates, adding to the interest of Block-PEC for cat-qubits.
Note that the described gain is not restricted to cat-qubits: Similar results can be obtained for Pauli-compatible circuits—i.e. Clifford circuits with Pauli controls under Pauli noise. This regime is explored in [27] that appeared after the completion of this work.
3.3 Computing Block-PEC coefficients